OpenStudy (anonymous):

would appreciate if someone can help me in solving below function limit lim {cos[(pi/6)+x] - sqrt3/2}/x ; when x approaching "0"

OpenStudy (fifciol):

\[\lim_{x \rightarrow 0}\frac{ \cos[ \frac{ \pi }{ 6}+x]-\frac{ \sqrt3 }{2 } }{ x }\] Use L'Hospital's rule

OpenStudy (fifciol):

if you substitute for x zero you have \[\left[\frac{ 0 }{ 0 } \right]\] so you can use that rule

OpenStudy (fifciol):

Do you know how to do that?

OpenStudy (anonymous):

or else expand cos(x+pi/6) and then solve

OpenStudy (fifciol):

yes, if you don't know what derivative is expand cos(x+pi/6)